# Dynamic thermalization on noisy quantum hardware

### v1

@@ Abstract
$\langle\langle{\cal O}(t>t_{\rm eq})\rangle\rangle={\rm Tr}\,\rho_{\rm th}(T(\varepsilon_{\rm in})){\cal O},$ (4)
traced with the thermal density matrix with the pre-quench Hamiltonian of the system. The temperature $T$ appears due to the quenched interaction and depends on the energy of the initial state $\varepsilon_{\rm in}$. The system thermalizes to a positive or negative temperature state as it has a bounded spectrum and no kinetic energy [9].
We experimentally examine our thermalization mechanism (3,4) in a simulation on an IBM Quantum computer (IBMQ) for a system of four qubits and read out the occupation probabilities of the many-body energy states upon equilibration. For the system’s initial state given by the ground state of the qubits, the average occupation probabilities relax to the Gibbs distribution. Starting from the highest excited state leads to the inversely populated state ($T<0$). As a result, both the observables and their fluctuations evaluated from the noisy quantum computer’s error-mitigated output display thermal behavior with positive or negative absolute temperatures.
Being implemented on noisy intermediate-scale quantum (NISQ) hardware, our approach allows probing thermalization in finite-size quantum systems in a highly controlled environment and may be utilized among nature-inspired quantum algorithms for thermal state preparation [10, 11]. Advantageously, it does not require long-time evolution or increasing the system size due to auxiliary qubits aiming to simulate a thermal bath.
The model.— Quenching a closed system is a natural way of probing its equilibration dynamics and subsequent thermalization [12]. For example, thermalization may occur via a non-linear classical mode and the development of turbulence, leading to eventual energy equipartition [13], or due to switching on a highly non-local interaction in a quantum system resulting in its equilibration to an infinite temperature state [14]. Taking this dynamical perspective, we organize our quench protocol as follows.
@@
describes the interaction between the qubits with the coupling constants $J_{j_{1}j_{2}}$ drawn from the Gaussian unitary ensemble with zero mean $\langle J_{j_{1}j_{2}}\rangle=0$ and finite variance $\langle|J_{j_{1}j_{2}}|^{2}\rangle=J^{2}/N$. Accordingly, one may understand the interaction term (6) as the Hamiltonian of random non-interacting fermions [15] projected onto the qubits’ basis via the Jordan-Wigner transform. When ${\cal V}$ is combined with $H_{0}$ in the full Hamiltonian (5), the latter describes an interacting theory from fermions’ or qubits’ perspective giving rise to the quantum many-body dynamics we address below.
@@
Having established equilibration (3), we explore the relaxation process further. Generically, describing the equilibrium state of a quantum many-body system requires a generalized Gibbs ensemble to account for the conserved quantities. The generalized Gibbs ensemble reduces to the conventional Gibbs ensemble if the conserved quantities are limited to total energy, particle number, and momentum [21]. Since our quench protocol respects only the total energy conservation on average, we compare the averaged observables upon equilibration with those predicted by the usual Gibbs distribution. For the comparison, we consider a Gibbs state $\rho_{\rm th}(T)=e^{-H_{0}/T}/({\rm Tr}\,e^{-H_{0}/T})$ with the initial Hamiltonian $H_{0}$. By doing that, we aim to achieve thermal behavior for the initial system, assuming that the emergent temperature $T$ absorbs the effects of the interaction. At $t=10/J>t_{\rm eq}$, we observe that while the occupation probabilities, $\langle n_{l}(t)\rangle=|\langle l|\psi(t)\rangle|^{2}$, stay irregular for each individual realization of the quench protocol, the average occupation probabilities (8) are well-approximated with the occupation probabilities $n_{l}^{\rm th}(T)={\rm Tr}\rho_{\rm th}(T)|l\rangle\langle l|$. To illustrate this agreement, we fit the average occupation probabilities (8) with the thermal ones obeying the Gibbs distribution
$n_{l}^{\rm th}(T)=\frac{e^{-\varepsilon_{l}/T}}{Z},\quad Z(T)=\sum\limits_{l=0}^{L-1}e^{-\varepsilon_{l}/T},$ (9)
with the temperature $T$ being a fitting parameter. We compare $\langle\langle n_{l}(t)\rangle\rangle$ with the thermal fit $n_{l}^{\rm th}(T)$ in Fig. 1d.
@@
$D(\langle\langle n_{l}(t)\rangle\rangle||n_{l}^{\rm th}(\overline{T}))\!=\!-S(t)\!+\!\frac{(\langle\langle H_{0}(t)\rangle\rangle\!-\!F(\overline{T}))}{\overline{T}},$ (10)
@@
To fulfill the first task, we adapt variational circuit recompilation [34, 35] to approximate the evolution operator $U(t)$ by a parametrized quantum circuit with $99.8\%$ accuracy, optimizing the tensor network decomposition of $U(t)$, before plugging it into the quantum computer (see Appendix A). Using this method, we reduce the number of CNOT gates required for the simulation to $N_{\rm CNOT}=60$, discarding the need for Trotterization, which enables one to reach arbitrary times.
We address the second task using an extension of the randomized compiling (RC) approach [36] converting the noise, not only on qubits to which a CNOT gate is applied but also on the neighboring qubits, to an incoherent noise channel (see Appendix B). The incoherent noise contains local and global components, where the latter acts uniformly on every qubit in the circuit. To account for the global incoherent noise, we approximate the noisy quantum computer’s output for average occupation probabilities (8) upon equilibration as a weighted mixture of the average occupations unaffected by the global incoherent noise and a uniform distribution describing a maximally mixed (infinite temperature) state. Aiming to see if the observables evaluated on the quantum computer display thermal behavior, we chose the average occupations unaffected by the global incoherent noise in the form of the finite temperature Gibbs distribution (9) so that the noisy average occupation probabilities read
@@
The temperature obtained from the quantum simulation is higher than the temperature from the thermal prediction of the exact dynamics, as we observe from the occupations’ slopes in Fig. 2b. We attribute this heat-up to the part of the hardware noise so far unattended. Indeed, earlier, we have discussed the effect of the global incoherent noise, yet the local incoherent noise may be present in the system. We account for the remaining noise using the zero-noise extrapolation (ZNE) method [37, 38, 39, 40, 41, 42], which involves repeating the experiment with artificially amplified noise and extrapolating the measurement results to the noiseless limit. For incoherent noise, a linear change in noise amplification generally leads to exponential change in an observable [39]. From this perspective, the inverse temperature $\beta=1/T$ in the exponent of thermal occupations in Eq. (11) may absorb the impact of the remaining noise. Elaborating on this observation, we repeat the experiment with the tripled noise by tripling every CNOT gate in the circuit. The average occupation probabilities in the quantum simulation with tripled noise amplification (QC3) remain well within the incoherent noise approximation, (11) as shown in Fig. 2a. In turn, the temperature determined from the QC3 data (see Fig. 2b) increases further compared to the QC result, highlighting how the hardware noise on NISQ devices can lead to the temperature rise in the simulated observables. In Fig. 2c, we extrapolate the inverse temperature to the noiseless case using a linear fit based on two noise amplification values (QC, QC3). As the inverse temperature appears in the exponent of the noisy occupations (11), our method effectively performs an exponential extrapolation on the observables after clearing the effect of the global incoherent noise from the quantum data. As a result, the value of $\beta_{\rm ZNE}=(1.29\pm 0.12)/J$ determined from ZNE lies within the error bar with $\beta_{\rm ED}=(1.21\pm 0.04)/J$ found from fitting the occupations computed in the exact dynamics with the Gibbs distribution (9).
@@
Discussion.— We show how thermal observables can dynamically emerge in a closed quantum system as a result of averaging over repeated instances of a global quench initializing Gaussian random non-local interactions between the system’s subparts. Averaging the observables over many realizations of the quench protocol leads to their equilibration (3) and thermalization (4) in time $t_{\rm eq}\sim{\cal O}(1)/J$, where $J$ is the characteristic interaction strength. The resulting averaged many-body occupation probabilities in the equilibrium regime, $t>t_{\rm eq}$, follow the Gibbs distribution with positive or negative absolute temperatures depending on the initial state’s energy, conserved on average in the course of after-the-quench dynamics. One can additionally tune the absolute value of equilibrium temperature by adjusting the interaction strength $J$. This dynamic thermalization is particularly convenient to examine on NISQ devices such as IBMQ since it does not rely on the long-time evolution, large system size, or auxiliary qubits to emulate a thermostat but solely on the repetition of the quench protocol with different sets of random coupling constants.
@@
Adapting the established error-mitigation techniques, we implement our thermalization protocol on the IBMQ machine using the circuit depth of $60$ CNOT gates. We obtain thermal occupations for the sixteen many-body energy states of the four-qubit system, populated according to positive or negative absolute temperatures depending on the choice of the initial state. Respectively, the observables and their fluctuations found in quantum simulation agree with the thermal predictions. This utility of the digital quantum simulator for producing thermal observables may be attributed to the notion of the quantum computer as an open system and our implementation of the thermalization protocol. First, we average the expectation values over realizations of the quench protocol aiming at mimicking the thermodynamic limit and facilitating thermalization in a relatively small system. Second, we apply zero-noise extrapolation that resembles taking a dissipationless limit in an open quantum system. Genuinely, these two limits are non-commutative, and their correct order, naturally implemented in our protocol, may be essential to justify the occurrence of the fluctuation-dissipation theorem in a quantum system [43]. Alongside, in our simulation, the time for the observables to equilibrate is ${\cal O}(1)/J$, similar to the timescale of irreversible dynamics $T_{3}$ in an open quantum system, bypassing which an unconstrained system thermalizes [44].
@@
The advantage of the considered quench protocol is its inherently dynamic nature, which allows studying the occurrence of thermalization or ergodicity breaking in quantum systems in real-time, efficiently engaging the limited resources of NISQ processors. We apply our quench protocol to study the emergence of the negative absolute temperature equilibrium on a digital quantum platform. While absolute temperatures are usually restricted to positive values due to kinetic energy, equilibrium at negative absolute temperatures is possible under certain conditions, in systems with a bounded energy spectrum. Bringing negative absolute temperature equilibrium states to reality inspired significant experimental efforts due to their non-trivial thermodynamic properties, e.g., losing entropy when absorbing energy. By now, negative temperature states were experimentally observed in a few physical systems such as nuclear spins [22, 23], cold atoms [24, 25, 28], two-dimensional quantum superfluids [26, 27], optical fibers [29], and a two-level system environment in superconducting qubits [30]. Here, we present equilibrium state with negative absolute temperature arising from the relaxation process in a quantum system realized digitally on a quantum computer, demonstrating the utility of the existing quantum processors for current research in statistical mechanics.
@@
Acknowledgments.— We thank Alexander Balatsky, Igor Gornyi, Archana Kamal, Alexander Mirlin, Horacio Pastawski, Anatoli Polkovnikov, and Alexander Shnirman for stimulating discussions. HP acknowledges support from the Horizon Europe program HORIZON-CL4-2021- DIGITALEMERGING-01-30 via the project 101070144 (EuRyQa). HP and TS acknowledge funding from the state of Baden-Württemberg through the Kompetenzzentrum Quantum Computing Project QC4BW. AIP was supported by the German Ministry of Education and Research (BMBF) within the project QSolid (FKZ: 13N16151). NVG was supported by the Department of Physics and Astronomy, Dartmouth College.
Appendix A Variational circuit recompilation

### v2

@@ Abstract
$\langle\langle{\cal O}(t>t_{\rm eq})\rangle\rangle={\cal O}_{\rm th}(T(\varepsilon_{\rm in})),$ (4)
where the temperature $T$ appears due to the quenched interaction and depends on the energy of the initial state $\varepsilon_{\rm in}$. The system thermalizes to a positive or negative temperature state as it has a bounded spectrum and no kinetic energy [9]. Below, we demonstrate this approach leads to thermal observables for the pre-quenched system without an external thermal bath.
We experimentally examine our thermalization mechanism (3,4) in a simulation on an IBM Quantum computer (IBMQ) for a system of four qubits. Upon equilibration, we read out the occupation probabilities of the many-body energy states of the system described by the pre-quench Hamiltonian. For the system’s initial state given by the ground state of the qubits, the average occupation probabilities relax to the Gibbs distribution. Starting from the highest excited state leads to the inversely populated state ($T<0$). As a result, both the observables and their fluctuations evaluated from the noisy quantum computer’s error-mitigated output display thermal behavior with positive or negative absolute temperatures.
Being implemented on noisy intermediate-scale quantum (NISQ) hardware, our approach allows probing thermalization in finite-size quantum systems in a highly controlled environment and may be utilized among nature-inspired quantum algorithms for thermal state preparation [10, 11]. Considerable experimental effort has been spent on sampling observables agreeing with the Gibbs state prediction on a digital (gate-based) quantum computer. Some recent experiments are based on a quantum imaginary time evolution algorithm [12, 13] or a variational approach with thermofield double states [14, 15], while another incorporates a thermal bath using ancilla qubits [11]. These experiments achieve thermalization for relatively small systems of $2$ to $4$ qubits. One concludes thermalization from the agreement of the observables measured at the end of the simulation with the Gibbs state prediction. Here, we also address thermalization in a small system on a digital quantum computer. We rely on the post-quench evolution and averaging the observables over realizations of the randomized auxiliary interaction. Our protocol is inherently dynamical, enabling studying the formation of thermal observables in real time. Furthermore, it relies on classical averaging of quantum observables corresponding to an individual realization of the quench protocol, unloading precious resources of the quantum simulator. Advantageously, it does not require long-time evolution or increasing the system size due to auxiliary qubits aiming to simulate a thermal bath.
[…]
The model.— Quenching a closed system is a natural way of probing its equilibration dynamics and subsequent thermalization [16]. For example, thermalization may occur via a non-linear classical mode and the development of turbulence, leading to eventual energy equipartition [17], or due to switching on a highly non-local interaction in a quantum system resulting in its equilibration to an infinite temperature state [18, 19]. Taking this dynamical perspective, we organize our quench protocol as follows.
@@
describes the interaction between the qubits with the coupling constants $J_{j_{1}j_{2}}$ drawn from the Gaussian unitary ensemble with zero mean $\langle J_{j_{1}j_{2}}\rangle=0$ and finite variance $\langle|J_{j_{1}j_{2}}|^{2}\rangle=J^{2}/N$. Accordingly, one may understand the interaction term (6) as the Hamiltonian of random non-interacting fermions [20] projected onto the qubits’ basis via the Jordan-Wigner transform. While both $H_{0}$ and ${\cal V}$ are integrable their combination is not. Indeed, when ${\cal V}$ is combined with $H_{0}$ in the full Hamiltonian (5), the latter is chaotic displaying Wigner-Dyson level statistics [21] as discussed in more detail in Appendix A. As such, the Hamiltonian (5) describes an interacting theory from fermions’ or qubits’ perspective giving rise to the quantum many-body dynamics we address below.
@@
Having established equilibration (3), we explore the relaxation process further. Since the Hamiltonian $H=H_{0}+{\cal V}$ is chaotic, it is natural to expect that the system equilibrates and thermalizes if it is large enough [6]. For a small system of $N=4$ qubits, a dynamic observable averaged over an ensemble of realizations equilibrates after passing an interaction-dependent timescale $t_{\rm eq}$, even though an observable computed with an individual realization of the quench protocol may severely oscillate. Generally, one can present the evolved state of the system as $|\psi(t)\rangle=e^{-i(H_{0}+{\cal V})t}|\psi_{\rm in}\rangle=\sum_{l}c_{l}(t)|l\rangle$, where we specifically choose the basis of eigenstates of the initial Hamiltonian $H_{0}$. Accordingly, the coefficients of this representation $c_{l}(t)=\langle l|\psi(t)\rangle$ are related to the probability of occupying the $l$-th state at time $t$: $\langle n_{l}(t)\rangle=|c_{l}(t)|^{2}$. For $t>0$, these occupation probabilities begin to deviate from their initial values and become dynamically redistributed, evolving with the chaotic Hamiltonian (5). If the system equilibrates, the probabilities to occupy the $l$-th state of $H_{0}$ should become stationary, describing an equilibrium probability distribution. Here we check if the average occupation probabilities become thermal.
Generically, describing the equilibrium state of a quantum many-body system requires a generalized Gibbs ensemble to account for the conserved quantities. The generalized Gibbs ensemble reduces to the conventional Gibbs ensemble if the conserved quantities are limited to total energy, particle number, and momentum [27]. Since our quench protocol respects only the total energy conservation on average, we compare the averaged occupation probabilities upon equilibration with those predicted by the usual Gibbs distribution:
$n_{l}^{\rm th}(T)=\frac{e^{-\varepsilon_{l}/T}}{Z},\quad Z(T)=\sum\limits_{l=0}^{L-1}e^{-\varepsilon_{l}/T}.$ (9)
By doing that, we aim to achieve thermal occupations for the initial system, assuming that the emergent temperature $T$ absorbs the effects of the interaction. At $t=10/J>t_{\rm eq}$, we observe that while the occupation probabilities $\langle n_{l}(t)\rangle$ stay irregular for each individual realization of the quench protocol, the average occupation probabilities (8) are well-approximated with the occupation probabilities $n_{l}^{\rm th}(T)$. To illustrate this agreement, we fit the average occupation probabilities (8) with the thermal ones (9) obeying the Gibbs distribution with the temperature $T$ being a fitting parameter. We compare $\langle\langle n_{l}(t)\rangle\rangle$ with the thermal fit $n_{l}^{\rm th}(T)$ in Fig. 1d.
@@
$D(\langle\langle n_{l}(t)\rangle\rangle||n_{l}^{\rm th}(\overline{T}))\!=\!-S(t)\!+\!\frac{\langle\langle H_{0}(t)\rangle\rangle\!-\!F(\overline{T})}{\overline{T}},$ (10)
@@
Figure 2: Thermalization dynamics of a four-qubit transverse Ising model with a periodic boundary condition in a paramagnetic phase after quenching ${\cal V}$. We use $\omega/g=4$ and $g=0.05J$. a) Expectation value of $H_{I}$ as a function of time after the quench. The gray curves show individual realizations of random couplings. The blue curve is the expectation value of $H_{I}$ averaged over $200$ realizations of random couplings. b) Probabilities to occupy eigenstates of $H_{I}$ read out upon equilibration. Alike in Fig. 1d, the blue squares show the average occupation probabilities (8) and red dots are the thermal fit (9). The spectrum of $H_{I}$ contains two pairs of degenerate states at $\varepsilon_{l}=\pm 0.4$ and a $4$-degenerate state at $\varepsilon_{l}=0$. c) Kullback-Leibler divergence between the average occupation probabilities and thermal occupations with the equilibrium temperature $\overline{T}\approx 0.87J$.
So far, we have achieved thermal occupations of the many-body energy states for the system described by the Hamiltonian $H_{0}$ of four non-interacting qubits, starting the protocol from an eigenstate of $H_{0}$. This setup can not evolve to a thermal equilibrium without the interaction potential ${\cal V}$ that breaks integrability and leads to eventual thermalization on average. To highlight the generality of our thermalization protocol, we choose an $N=4$ transverse Ising model with a periodic boundary condition, $H_{I}=-\omega\sum_{j=1}^{N}\sigma_{j}^{x}-g\sum_{j=1}^{N-1}\sigma_{j}^{z}\sigma_{j+1}^{z}-g\sigma_{N}^{z}\sigma_{1}^{z}$, as an initial Hamiltonian. This system is also integrable, yet the induced interaction (6) again leads to average thermal occupations of the eigenstates of $H_{I}$ with eigenenergies $\varepsilon_{l}$, as shown in Fig. 2. We tune the initial Hamiltonian in to a paramagnetic phase ($\omega=0.2J$, $g=0.05J$) and begin the evolution with a product state, $|\psi_{\rm in}\rangle=(1/\sqrt{2}\begin{pmatrix}1&1\end{pmatrix}^{\!\rm T})^{\otimes N}$, which is no longer an exact eigenstate of $H_{I}$: $\varepsilon_{\rm in}\neq\varepsilon_{l}$. Running $200$ realizations of random couplings, we observe in Fig. 2a that an average expectation value of $H_{I}$ equilibrates at $t\sim{\cal O}(1)/J$, where $J$ is the largest energy scale in the problem. We plot the occupation probabilities for $t=25/J$ in the equilibrium regime in Fig. 2b and find that mean occupation probabilities are well approximated by the Gibbs distribution. Note that the spectrum of $H_{I}$ has multiple degeneracies. Finally, in Fig. 2c, we show the Kullback-Leibler divergence (10) between the mean dynamic occupations and Gibbs distribution with an equilibrium temperature $\overline{T}$ determined from fitting the dynamic data averaged over an interval from $t=10/J$ to $t=50/J$ by thermal ones (9). These results are qualitatively similar to our previous analysis, namely, Figs. 1a,d,e.
@@
To fulfill the first task, we adapt variational circuit recompilation [40, 41] to approximate the evolution operator $U(t)$ by a parametrized quantum circuit with $99.8\%$ accuracy, optimizing the tensor network decomposition of $U(t)$, before plugging it into the quantum computer (see Appendix B). Using this method, we reduce the number of CNOT gates required for the simulation to $N_{\rm CNOT}=60$, discarding the need for Trotterization, which enables one to reach arbitrary times.
@@
We address the second task using an extension of the randomized compiling (RC) approach [42] converting the noise, not only on qubits to which a CNOT gate is applied but also on the neighboring qubits, to an incoherent noise channel (see Appendix C). The incoherent noise contains local and global components, where the latter acts uniformly on every qubit in the circuit. To account for the global incoherent noise, we approximate the noisy quantum computer’s output for average occupation probabilities (8) upon equilibration as a weighted mixture of the average occupations unaffected by the global incoherent noise and a uniform distribution describing a maximally mixed (infinite temperature) state. Aiming to see if the observables evaluated on the quantum computer display thermal behavior, we chose the average occupations unaffected by the global incoherent noise in the form of the finite temperature Gibbs distribution (9) so that the noisy average occupation probabilities read
@@
The temperature obtained from the quantum simulation is higher than the temperature from the thermal prediction of the exact dynamics, as we observe from the occupations’ slopes in Fig. 3b. We attribute this heat-up to the part of the hardware noise so far unattended. Indeed, earlier, we have discussed the effect of the global incoherent noise, yet the local incoherent noise may be present in the system. We account for the remaining noise using the zero-noise extrapolation (ZNE) method [43, 44, 45, 46, 47, 48], which involves repeating the experiment with artificially amplified noise and extrapolating the measurement results to the noiseless limit. For incoherent noise, a linear change in noise amplification generally leads to exponential change in an observable [45]. From this perspective, the inverse temperature $\beta=1/T$ in the exponent of thermal occupations in Eq. (11) may absorb the impact of the remaining noise. Elaborating on this observation, we repeat the experiment with the tripled noise by tripling every CNOT gate in the circuit. The average occupation probabilities in the quantum simulation with tripled noise amplification (QC3) remain well within the incoherent noise approximation, (11) as shown in Fig. 3a. In turn, the temperature determined from the QC3 data (see Fig. 3b) increases further compared to the QC result, highlighting how the hardware noise on NISQ devices can lead to the temperature rise in the simulated observables. In Fig. 3c, we extrapolate the inverse temperature to the noiseless case using a linear fit based on two noise amplification values (QC, QC3). As the inverse temperature appears in the exponent of the noisy occupations (11), our method effectively performs an exponential extrapolation on the observables after clearing the effect of the global incoherent noise from the quantum data. As a result, the value of $\beta_{\rm ZNE}=(1.29\pm 0.12)/J$ determined from ZNE lies within the error bar with $\beta_{\rm ED}=(1.21\pm 0.04)/J$ found from fitting the occupations computed in the exact dynamics with the Gibbs distribution (9). The canonical prediction for $\beta_{c}=1/T_{c}=1.20/J$ based on energy conservation well agrees with exact dynamics and experimental value of the inverse temperature upon error mitigation.
@@
Discussion.— We show how thermal observables can dynamically emerge in a closed quantum system as a result of averaging over repeated instances of a global quench initializing Gaussian random non-local interactions between the system’s subparts. After the quench, the system pursues evolution governed by the chaotic Hamiltonian falling into the class of Gaussian unitary ensemble of random matrices. Averaging the observables over many realizations of the quench protocol leads to their equilibration (3) and thermalization (4) in time $t_{\rm eq}\sim{\cal O}(1)/J$, where $J$ is the characteristic interaction strength. The resulting averaged probabilities of occupying eigenstates of the initial system in the equilibrium regime, $t>t_{\rm eq}$, follow the Gibbs distribution with positive or negative absolute temperatures depending on the initial state’s energy, conserved on average in the course of after-the-quench dynamics. One can additionally tune the absolute value of equilibrium temperature by adjusting the interaction strength $J$. This dynamic thermalization is particularly convenient to examine on NISQ devices such as IBMQ since it does not rely on the long-time evolution, large system size, or auxiliary qubits to emulate a thermostat but solely on the repetition of the quench protocol with different sets of random coupling constants.
@@
Adapting the established error-mitigation techniques, we implement our thermalization protocol on the IBMQ machine using the circuit depth of $60$ CNOT gates. We obtain thermal occupations for the sixteen many-body energy states of the four-qubit system, populated according to positive or negative absolute temperatures depending on the choice of the initial state. Respectively, the observables and their fluctuations found in quantum simulation agree with the thermal predictions. This utility of the digital quantum simulator for producing thermal observables may be attributed to the notion of the quantum computer as an open system and our implementation of the thermalization protocol. First, we average the expectation values over realizations of the quench protocol aiming at mimicking the thermodynamic limit and facilitating thermalization in a relatively small system. One way to think of thermalization in a small system is to consider it as a subpart of a much bigger system in the thermodynamic limit, with the rest of the system acting as a thermal bath. A recent observation notes that averaging the evolution of a system with randomized interaction results in effective dissipative dynamics [19]. Another recent study shows that adding a disorder in a simulated system in analog quantum simulations substantially reduces the necessary size of a bath for infinite temperature thermalization [49]. In our experiment, we start with a pure state and observe dynamical relaxation to a finite temperature equilibrium for averaged observables without involving a thermal bath, with the mitigated noise seemingly falling into the same effectively dissipative channel and raising the temperature of the simulated observables. Second, we apply zero-noise extrapolation that resembles taking a dissipationless limit in an open quantum system. Genuinely, these two limits are non-commutative, and their correct order, naturally implemented in our protocol, may be essential to justify the occurrence of the fluctuation-dissipation theorem in a quantum system [50]. Alongside, in our simulation, the time for the observables to equilibrate is ${\cal O}(1)/J$, similar to the timescale of irreversible dynamics $T_{3}$ in an open quantum system, bypassing which an unconstrained system thermalizes [51].
@@
The advantage of the considered quench protocol is its inherently dynamic nature, which allows studying the occurrence of thermalization or ergodicity breaking in quantum systems in real-time, efficiently engaging the limited resources of NISQ processors. We apply our quench protocol to study the emergence of the negative absolute temperature equilibrium on a digital quantum platform. While absolute temperatures are usually restricted to positive values due to kinetic energy, equilibrium at negative absolute temperatures is possible under certain conditions, in systems with a bounded energy spectrum. Bringing negative absolute temperature equilibrium states to reality inspired significant experimental efforts due to their non-trivial thermodynamic properties, e.g., losing entropy when absorbing energy. By now, negative temperature states were experimentally observed in a few physical systems such as nuclear spins [28, 29], cold atoms [30, 31, 34], two-dimensional quantum superfluids [32, 33], optical fibers [35], and a two-level system environment in superconducting qubits [36]. Here, we present an equilibrium state with negative absolute temperature arising from the relaxation process in a quantum system realized digitally on a quantum computer, demonstrating the utility of the existing quantum processors for current research in statistical mechanics.
@@
Acknowledgments.— We thank Alexander Balatsky, Igor Gornyi, Archana Kamal, Alexander Mirlin, Horacio Pastawski, Anatoli Polkovnikov, and Alexander Shnirman for stimulating discussions. HP acknowledges support from the Horizon Europe program HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via the project 101070144 (EuRyQa). HP and TS acknowledge funding from the state of Baden-Württemberg through the Kompetenzzentrum Quantum Computing Project QC4BW. AIP was supported by the German Ministry of Education and Research (BMBF) within the project QSolid (FKZ: 13N16151). NVG was supported by the Department of Physics and Astronomy, Dartmouth College.
Figure 5: Probability distribution of the ratio of consecutive level spacings for the Hamiltonian (5) in the main text. The blue histogram shows the probability distribution for a system of $N=4$ qubits with the same $M=100$ realizations we use for exact and quantum dynamics. We display the data averaged over $12$ ratios of consequential level spacings without two pairs of states at the edges of the spectrum. The red histogram is done for the system with $N=10$ qubits over $M=1000$ realizations, averaging the data over $\sim 10\%$ out of $2^{10}$ of many-body energy levels. For comparison, the black dashed line shows the GUE predicted level statistics. The gray dotted line displays Poisson (integrable) behavior put for a visual reference. Both ($N=4$, $M=100$) and ($N=10$, $M=1000$) datasets saturate the GUE curve and possess the GUE characteristic value of the mean adjacent gap ratio $\langle r\rangle_{\rm GUE}\approx 0.6$.
Figure 6: Equilibration of the multi-qubit system in the initial state $|\psi_{\rm in}\rangle=|\downarrow\downarrow\ldots\downarrow\rangle$ after the quench. We plot the expectation value of $H_{0}$ as a function of time for individual realizations of random couplings (gray curves), averaged over $M=100$ realizations (solid blue curve), averaged over $M=10$ realizations (dash-dotted green curve), and averaged over $M=20$ realizations (dashed red curve) for different system sizes: a) for $N=4$ qubits, b) for $N=6$ qubits, and c) $N=8$ qubits. The vertical black line marks the equilibration timescale at $t=2.5/J$.
Appendix A Choice of thermalization protocol
The choice of the interaction is essential for our protocol so that there would be no constraints preventing thermalization after the quench. We began with a “constrained” non-interacting Hamiltonian $H_{0}$ that can not be the sole source of thermalization dynamics. In our case, thermalization occurs due to the quenched random interaction ${\cal V}$ (with zero mean strength) between the qubits. Owing to the zero mean strength interaction, adding the latter to the initial Hamiltonian conserves only the total energy equal to the initial energy of the system on average. The inter-qubit interaction $\cal V$ maps onto the random free fermions, also known as the SYK2 model, an explicitly non-interacting theory. However, the combined Hamiltonian $H=H_{0}+{\cal V}$ is non-integrable, as may be seen from its level statistics.
The notion of energy level statistics allows characterizing integrability, localization, and chaos in quantum many-body systems. Chaotic behavior, closely related to thermalization [6], manifests in statistical properties of a system’s many-body energy levels, falling in the Wigner-Dyson statistics described by a Gaussian orthogonal (GOE), unitary (GUE), or symplectic (GSE) ensemble of random matrices [52]. On the contrary, integrability and localization are associated with level statistics following a Poisson distribution [53]. Accordingly, changing disorder in a system may lead to a localization transition from Wigner-Dyson to Poisson level statistics [54, 55, 56, 57, 58, 59].
To quantify the level statistics in our model, we use the probability distribution for a ratio $0\leq r_{l}\leq 1$ determined from two consecutive level spacings in a given part of a many-body spectrum [54]: $r_{l}={\rm min}\{\delta_{l+1},\delta_{l}\}/{\rm max}\{\delta_{l+1},\delta_{l}\}$, where $\delta_{l}=\tilde{\varepsilon}_{l+1}-\tilde{\varepsilon}_{l}$. Note that here $\tilde{\varepsilon}_{l}$ are the eigenvalues of $H$. For the adjacent gap ratio $r_{l}$ following Wigner-Dyson statistics, the probability distribution is $P(r)=2(1+r)^{\beta}/(1+r+r^{2})^{1+\frac{3\beta}{2}}/Z_{\beta}$ with $Z_{\beta}$ being a normalization constant and $\beta$ is a symmetry index (do not confuse with an inverse temperature) equal to $1,2$, and $4$ for GOE, GUE, and GSE, predicting mean values of the gap ratio $\langle r\rangle_{\rm GOE}\approx 0.54$, $\langle r\rangle_{\rm GUE}\approx 0.6$, and $\langle r\rangle_{\rm GSE}\approx 0.68$ [56]. In the integrable case, for the Poisson spectrum, the distribution of the gap ratio is $P(r)=2/(1+r)^{2}$ with $\langle r\rangle_{\rm Poisson}\approx 0.39$ [54]. Following earlier studies [54, 55, 56, 57, 58, 59], we consider our total Hamiltonian $H=H_{0}+{\cal V}$ in the regime when the system is large enough ($N=10$ with $2^{N}$ many-body energy levels) and plot the histogram for the gap ratio distribution for $M=1000$ disorder realizations averaged over $\sim 10\%$ of levels in Fig. 6. The distribution function is in complete agreement with GUE prediction and reproduces the mean adjacent gap ratio $\langle r\rangle_{\rm GUE}\approx 0.6$. Reducing the system’s size and the number of realizations to what we use in our experiment and choosing the parameters accordingly, we observe that the level statistics of the model remain well-described by GUE in Fig. 6, highlighting an absence of integrability or localization.
Having established that the model is chaotic, i.e., described by a random matrix theory, one expects the observables evolved with the Hamiltonian $H$ to equilibrate and thermalize. However, one realization of random couplings may not be enough for equilibration. In Fig. 6, we show the after-the-quench equilibration process for $\langle H_{0}(t)\rangle$ for different system sizes ($N=4,6,8$), averaging the expectation values with several numbers of realizations $M=10,20,100$. Increasing the number of realizations leads to less finite-size oscillations in the average observable in the equilibrium regime. The larger the system, the fewer disorder realizations are needed to suppress the remaining oscillations in the observable. In turn, the time the system evolves to achieve average stationary observables $t_{\rm eq}\sim{\cal O}(1)/J$ does not depend on the system’s size, being reminiscent of a characteristic timescale of irreversible dynamics $T_{3}$ in open quantum systems, solely defined by an inverse interaction amplitude, bypassing which the system relaxes toward thermal equilibrium [51]. This resemblance with open quantum systems is well-aligned with a recent finding that the average time evolution by a random interacting Hamiltonian is equivalent to dissipative dynamics [19]. A complementary observation has been made in the recent work on deep thermalization at infinite temperature tailored for analog quantum simulators, noting that adding a random potential in a simulated Hamiltonian reduces the required amount of the bath qubits [49].
Equilibration in the system results in thermal observables. In particular, we use the probabilities to occupy the many-body energy eigenstates of the initial Hamiltonian $H_{0}$. After the quench, the average occupation probabilities relax to a finite temperature Gibbs distribution. This protocol allows dynamically obtaining equilibrium observables at finite temperatures for a given system on a digital quantum processor without simulating a thermal bath.
Appendix B Variational circuit recompilation